OpenStudy (anonymous):

Solve the triangle. A = 51°, b = 14, c = 6

OpenStudy (jack1):



OpenStudy (jack1):

u can use the sine rule here: a/sin A = b/sin B = c/sin C \[\frac{ a }{ \sin A }= \frac{ b }{ \sin B }= \frac{ c }{ \sin C }\]

OpenStudy (jack1):

or u can use the cosine rule: \[c^2 = a^2 + b^2 - 2ab \times \cos(C)\] rearrange as you currently know A, and solve for c... etc

OpenStudy (jack1):

so which will u use @Loveiskey18 ?

OpenStudy (anonymous):

The second formula.

OpenStudy (jack1):

right, first off, and what did u get for A?

OpenStudy (jack1):

dammit, sorry, little "a"?

OpenStudy (anonymous):

2601.

OpenStudy (anonymous):

nevermind, disregard that

OpenStudy (anonymous):

You know what, let's use the first formula, I thought that the second one was going to be simple, that is why I chose it...

OpenStudy (jack1):

will be difficult until u know both one side and it's corresponding angle, otherwise you'll have to do some funky subsitiution... save the sine rule for after you solve "a" using cosine rule

OpenStudy (jack1):

so a^2 = b^2 + c^2 -2*b*c* cos (A)

OpenStudy (jack1):

a^2 = b^2 + c^2 -2*b*c* cos (A) a^2 = 14^2 + 6^2 -2*14*6* cos (51) ... so times it all out and solve, u get a^2 = 126.27 so a = 11.2372 ...ish now use the sine rule from here, yeah?

OpenStudy (anonymous):

yes

OpenStudy (jack1):

sweet, good luck! did the above all make sense tho?

OpenStudy (anonymous):

Actually it did. Thank You, I really do appreciate your help. :)